OpenStudy (anonymous):

Find the exact value of each logarithm. log1/2 16 can someone show me the steps

OpenStudy (anonymous):

\[\left(\frac{1}{2}\right)^x=16\] solve for \(x\)

OpenStudy (anonymous):

\[\log_\frac{1}{2} 16 \]\[=\frac{\log 16}{\log\frac{1}{2}}\]\[\log 16 = \log (2^4)\]\[\log \frac{1}{2} = \log (2^{-1})\] Using \[\log m^n = n\log m\] You can express log16 and log(1/2) in terms of log2. Then, it should not be difficult.

OpenStudy (radar):

\[\log 1^{-x}/(1/2)=16\]

OpenStudy (radar):

\[\log 2^{-x}=16\]\[2^{-x}=16\]\[-x=4\] x=-4 For some reason, I feel I have violated some mathematical rule in this approach. Please tell me if I erred.